# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

### v1

@@ Abstract
LABEL:fig:artery2_images shows the vascular structure imaged by rotational angiography of the right carotid artery for the second aneurysm case, in addition to the surface mesh (utilized in this study) obtained from a segmentation of the region of interest. LABEL:fig:artery2_streamlines shows streamline data from a CFD simulation for this case (executed outside of this study [43]). The velocity at each point along these streamlines is used to construct a metric tensor field to guide adaptivity for anisotropic mesh generation. A velocity gradient tensor is calculated at each point which can be mapped to a 3x3 positive definite matrix. The magnitude (eigenvalue) of each velocity gradient tensor is used to apply a stretching factor to the positive definite matrix which ensures that elements are stretched in the dominant direction of the velocity. An isotropic volume mesh is first generated using the surface mesh of the second case. A k-nn search is utilized to find the 3 closest velocity points to each mesh volume vertex (which are then used to calculate the velocity gradient tensor for each mesh point). OpenMP is again utilized to parallelize the processing of the mesh points. The isotropic volume mesh is processed as a background mesh with the velocity-based metric tensor field to generate anisotropic meshes. Both the anisotropic and isotropic meshes generated by CDT3D for the second case are evaluated.
5.2 Quantitative and Qualitative Analysis
@@
The results obtained from CDT3D are also compared to TetGen [61, 5] (replacing the adaptive anisotropic method with this adaptive isotropic method in the pipeline). TetGen is a sequential, Delaunay-based method that generates isotropic tetrahedral meshes. In addition to an anisotropic metric field, an isotropic sizing function is also constructed (based on the aforementioned approach in [21]) specifically for TetGen. We compare results between meshes generated by TetGen when it utilizes this sizing function and when it does not (tetrahedralization of the CBC3D surface). Additionally, we compare the TetGen isotropic meshes to isotropic meshes generated by CDT3D (again simply processing the same surface as a PLC - .stl).
@@
Table 1 shows the parameters used for each mesh generation method in the pipeline (except CBC3D). The first aneurysm case was also used to evaluate CBC3D in [27] (as it was compared to other isotropic I2M conversion methods from industry and academia). The same parameters utilized for CBC3D in that study are utilized in this evaluation. A wide range of values were tested for the value of k used in the k-nn search when building the sizing function and metric tensor fields for the first case. The optimal parameters were determined based on the best quality mesh that was generated. If certain parameters are not listed in table 1, then their default value was utilized. The optimal value for k with the isotropic runs of TetGen is 30. The optimal value for k in the metric-adapted anisotropic runs of CDT3D is 40. The CDT3D parameters nqual and nsmth were both increased to 10 while cdfn was changed to 0.8 to improve the final anisotropic mesh quality for the second aneurysm case (with no surface adaptation by CDT3D, explained later).
@@
TetGen (iso-sizing) r reconstruct mesh
@@
CBC3D takes approximately 61.59 seconds (1 min) to convert the segmented image of the first aneurysm into a mesh of approximately 272K tetrahedra and 63K points. LABEL:fig:Mesh_Cuts_artery1 shows cross sections of the volume meshes generated by each method, merged with the conforming AFLR boundary layer mesh. LABEL:fig:Mesh_Cuts_artery2 shows cross sections of the volume meshes generated for the second aneurysm case. Table 2 shows the number of elements generated by each method when processing the CBC3D mesh of aneurysm 1 and the surface mesh of aneurysm 2. The TetGen meshes contain fewer elements compared to the CDT3D meshes in both cases. Table 3 and table 4 show the runtime of each method for aneurysm cases 1 and 2, respectively. They also show the mesh generation/adaptation rates of each method (measured in elements generated per second). TetGen exhbits good performance when generating its small-size meshes. Although it is not listed in table 3, it should be noted that CDT3D exhibits good scalability, as its runtime is reduced to approximately 1 minute and its anisotropic adaptation rate becomes 27K when utilizing 40 cores. When utilizing 10 cores, CDT3D’s mesh generation/adaptation rate exceeds that of TetGen’s. For the second case, CDT3D completes anisotropic adaptation in about 49 seconds when utilizing 40 cores.
Table 2: Number of tetrahedra and points generated by each method when processing the CBC3D mesh converted from the segmented image of the first aneurysm and the surface mesh of the second aneurysm. ‘Tets’ is short for tetrahedra. K means thousand and M means million.
@@
TetGen (iso-sizing) Tets 120K -
@@
AFLR (bl) Tets 805K 105K
Points 144K 18K
@@
Table 3: The time spent (approximately in minutes) and the number of elements generated per second by each method (in parentheses) for small-size meshes is shown when processing the CBC3D mesh converted from the segmented image of the first aneurysm. Instead of elements per second, the values in parentheses for Build_Sizing/Metric represent the CBC3D volume mesh points processed per second. K means thousand.
@@
TetGen (iso-sizing) 0.1 (20K) - -
TetGen (iso) 0.06 (32K) - -
AFLR (bl) 0.2 (67K) - -
@@
Table 4: The time spent (approximately in seconds) and the number of elements generated per second by each method (in parentheses) for small-size meshes is shown when processing the surface mesh of the second aneurysm. K means thousand.
@@
Mesh fidelity is evaluated using a two-sided Hausdorff Distance (HD) metric (based on an open source implementation [23]): $\text{HD}=\max\{\text{HD}_{\text{I}\rightarrow\text{M}},\text{HD}_{\text{M}\rightarrow\text{I}}\}$, where $\text{HD}_{\text{I}\rightarrow\text{M}}$ is the value of the metric from the image to the mesh, and $\text{HD}_{\text{M}\rightarrow\text{I}}$ is the value of the metric from the mesh to the image. A low HD error indicates a high fidelity. Hausdorff Distance is computed between two point sets. The first point set contains the vertices located on the mesh surface and the second point set contains the voxels located on the boundaries of the segmented material. For the first aneurysm case (in LABEL:fig:artery1), $\text{HD}_{\text{I}\rightarrow\text{M}}$ is 4.51 while $\text{HD}_{\text{M}\rightarrow\text{I}}$ is 1.74. The HD is therefore 4.51. The fidelity of the mesh generated by CBC3D was also evaluated for the first aneurysm case in [27]. CBC3D was compared to several highly used I2M methods, exhibiting good fidelity while maintaining better smoothness compared to the other methods’ meshes. Surface refinement/adaptation is turned off for all mesh generation methods tested in this pipeline (except CBC3D). This ensures that the high fidelity and smooth surface generated by CBC3D remains intact.
For the isotropic meshes, qualitative results are examined with respect to element dihedral angles. LABEL:fig:artery1_dihedrals compares the quality of the isotropic meshes generated by each method for case 1 and LABEL:fig:artery2_dihedrals compares those generated for case 2. Table 5 shows the smallest and largest dihedral angles found in the isotropic meshes generated by each method. The CDT3D meshes contain better quality compared to the TetGen meshes. This is particularly seen in (b) and (c) of LABEL:fig:artery1_dihedrals and LABEL:fig:artery2_dihedrals, which show that the TetGen meshes contain more elements with small and large angles. For the adapted anisotropic meshes, qualitative results are examined with respect to their metric conformity. The goal of metric conformity is to create a unit grid, where edges are unit-length and elements are unit-volume with respect to the target metric. For calculating edge length, we adopted the same definition that appears in [7]. For two vertices a and b, an edge length in the metric $L_{e}$ can be evaluated using:
(5.1) $\begin{split}&L_{e}=\begin{cases}\frac{L_{a}-L_{b}}{log(L_{a}/L_{b})}&|L_{a}-L_{b}|>0.001\\ \frac{L_{a}+L_{b}}{2}&otherwise\end{cases}\\ &L_{a}=(v_{e}^{T}M_{a}v_{e})^{\frac{1}{2}},L_{b}=(v_{e}^{T}M_{b}v_{e})^{\frac{1}{2}}\end{split}$
and an element mean ratio shape measure can be approximated in the discrete metric as:
(5.2) $Q_{k}=\frac{36}{3^{1/3}}\frac{\left(|k|\sqrt{det(M_{mean})}\right)^{\frac{2}{3}}}{\sum_{e\epsilon L}v_{e}^{T}M_{mean}v_{e}},$

### v3

@@ 4.1 I2M Pipeline
For our second aneurysm test case, we utilize streamline data of the velocity field from a CFD simulation executed outside of this study [CFDDrivenTopology2024] (seen in Figure 5(b) of section 5). The velocity at each point along these streamlines is used to construct a metric tensor field to guide adaptivity for anisotropic mesh generation. Because this aneurysm case was only provided to us as a surface mesh, an isotropic volume mesh is first generated using the surface mesh. A k-nn search is utilized to find the 3 closest velocity points to each mesh volume vertex. Using the 3 closest points, a velocity gradient tensor is calculated at each mesh vertex which can be mapped to a 3x3 positive definite matrix. The magnitude (eigenvalue) of each velocity gradient tensor is used to apply a stretching factor to the positive definite matrix which ensures that elements are stretched in the dominant direction of the velocity. OpenMP is again utilized to parallelize the processing of the mesh points. The isotropic volume mesh is processed as a background mesh with the velocity-based metric tensor field to generate anisotropic meshes. A qualitative and performance evaluation of the pipeline is given in section 5.
4.2 Hierarchical Load Balancing
A hierarchical load balancing technique was previously seen to enhance the performance of the I2M Delaunay-based multi-threaded software PODM, which also utilizes a speculative execution model [foteinos2014high] (discussed in more detail in section 4.4). Consequently, we introduce a cc-NUMA-based hierarchical load balancing technique for the adaptive anisotropic parallel mesh generation software CDT3D. Before adaptation, CDT3D organizes tetrahedra into ”buckets” that are all assigned to pinned threads (i.e., each thread is pinned to a CPU core) [EwCCDT3D, DrakopoulosCDT3D]. When a thread finishes processing its list of buckets during adaptation, it is inserted into an idle list. Originally, after processing a bucket, a busy thread (those that still have buckets to process) would give a fraction of its buckets to the first thread in the idle list (if the list is not empty) to balance workloads. CDT3D has been updated so that a busy thread will now search the list of idle threads to find the pinned thread that is located within the shortest numa node distance among all the idle threads. cc-NUMA-based shared memory architectures organize CPU cores into numa nodes, where one numa node will consist of multiple CPU cores. When a thread creates new elements, their respective data structures are allocated in memory that is local to the numa node within which that pinned thread’s CPU core is located. Any CPU core (and its pinned thread) within that same numa node will relatively have the same memory access time to those data structures. Remote memory accesses (by CPU cores located in other numa nodes on the shared memory system) require greater access time. Consequently, it is ideal for busy threads to give workloads to threads that are pinned to CPU cores within the shortest numa node distance to reduce the number of remote memory accesses and limit the latency of such accesses during adaptation. Both the syscall [^1] https://man7.org/linux/man-pages/man2/syscall.2.html and numa [^2] https://man7.org/linux/man-pages/man3/numa.3.html libraries are utilized to get the numa node id of a CPU core and the distance between two numa nodes (given their ids), respectively. Algorithm 1 gives an overview of CDT3D’s hierarchical load balancing. We evaluate the effectiveness of such a technique compared to CDT3D’s original load balancing method in section 5.3.
Algorithm 1 CDT3D’s Hierarchical Load Balancing
NUMA_LOAD_BALANCE($threads_{1}...threads_{N}$, $buckets_{1}...buckets_{N}$)
Input: $threads_{1}...threads_{N}$ are the threads
$buckets_{1}...buckets_{N}$ is the list of buckets per thread in $threads_{1}...threads_{N}$
Output: Parallel adapted mesh
1:idle_thread_list = empty list
2:foreach $thread_{i}\in threads_{1}...threads_{N}$ do
3:  PIN_THREAD($thread_{i}$)
4:  $numa\_node\_id_{thread_{i}}$ = GET_NUMA_NODE_ID($thread_{i}$)
5:end for
6:foreach $thread_{i}\in threads_{1}...threads_{N}$ do $\triangleright$ Executed in parallel
7:  while $buckets_{i}$ is NOT empty do
8:   if idle_thread_list is NOT empty then
9:    closest_thread = idle thread w/ minimum numa node distance from $numa\_node\_id_{thread_{i}}$
10:    Assign fraction of $thread_{i}$’s buckets to closest_thread and remove them from $buckets_{i}$
11:    WAKE_UP(closest_thread)
12:   end if
13:   bucket = first bucket in $buckets_{i}$
14:   APPLY_MESH_OPERATIONS(bucket)
15:   Remove bucket from $buckets_{i}$
16:  end while
17:  idle_thread_list-¿insert($thread_{i}$)
18:  SLEEP($thread_{i}$)
19:end for
20:TERMINATE
4.3 Optimized Local Reconnection
Upon further study, we have found that a percentage of the elements within meshes generated by CDT3D may not fully satisfy qualitative criteria required by the target metric field. Despite this phenomena, CDT3D’s meshes have been compared to those generated by several state-of-the-art methods, exhibiting comparable overall quality [TsolakisEvaluation] and excellent performance when integrated into an adaptive CFD simulation pipeline [EwCCDT3D]. An example of this phenomena is that approximately 6-8% of elements generated for our second aneurysm case (where its metric field is scaled to 50 million complexity to generate about 100 million elements, seen in section 5) do not satisfy qualitative criteria by the end of execution. Furthermore, we found that the local reconnection operation wastes a significant amount of time attempting to improve these elements. The local reconnection operation is utilized within both phases of CDT3D (mesh adaptation and quality improvement), consisting of topological transformations (also called flips) [DrakopoulosCDT3D, EwCCDT3D]. Operations within each phase are executed repeatedly over numerous iterations until little to no new points are inserted (for mesh adaptation) or a user-defined threshold is reached. If CDT3D detects that elements do not satisfy qualitative criteria, then it will utilize the local reconnection operation to attempt to improve those elements. A problem is that it will attempt to improve these elements during every iteration, regardless of whether or not that element or its neighbors were modified in a previous iteration. This means that CDT3D will waste time attempting to improve the same elements repeatedly over numerous iterations. We found that on average, the local reconnection operation simply repeats itself (for more than one iteration) attempting to improve approximately 80% of the elements (the 6-8%) that do not satisfy qualitative criteria by the end of execution.
The local reconnection operation has now been optimized to avoid unnecessary repetition when attempting to improve element quality. It keeps track of whether or not flips have already been attempted for an element, and recognizes that it should not re-attempt improvement if the element or its neighbors have not been modified since the last attempt (from a previous iteration). The method also recognizes if flips should be re-attempted due to unsuccessful dependency locking (when attempting to improve the element during a previous iteration) under CDT3D’s speculative execution model. Algorithm 2 shows this new functionality. The amount of time saved and the impact of this optimization can be seen in section 5.4.
Algorithm 2 Optimized Local Reconnection
LOCAL_RECONNECTION(tetrahedra, current_iteration)
Input: tetrahedra is the set of tetrahedra upon which to attempt flips
current_iteration is the iteration number in which this operation is being executed (among the iterations for either mesh adaptation or quality improvement)
Output: tetrahedra with potentially improved quality
1:foreach tet $\in$ tetrahedra do
2:  neighborElements = set of tetrahedra that share a face with tet
3:  if (current_iteration $\neq$ 1 AND tet has NOT encountered unsuccessful dependency locking) then
4:   mainModified = false
5:   neighborsModified = false
6:   if (tet was created during current_iteration OR
7:    tet was modified by another operation during current_iteration) then
8:    mainModified = true
9:   end if
10:   if NOT mainModified then
11:    foreach neighbor $\in$ neighborElements do
12:      if (neighbor was created during current_iteration OR
13:    neighbor was modified by another operation during current_iteration) then
14:        neighborsModified = true
15:      end if
16:    end for
17:   end if
18:   if (NOT mainModified AND NOT neighborsModified) then
19:    Skip tet and continue to next iteration of for loop (line 1)
20:   end if
21:  end if
22:  if neighborElements are successfully locked then
23:   ATTEMPT_FLIPS(tet, neighborElements)
24:   Unmark tet for unsuccessful dependency locking
25:  else
26:   Mark tet with unsuccessful dependency locking
27:  end if
28:end for
4.4 Function Approximation for Medical Imaging
Although it may be inadequate in resolving features of some stent geometries and its output mesh quality is sometimes inferior to that of CBC3D’s [EwCCBC3D], we gauge the feasibility of utilizing the image-to-mesh conversion software PODM [foteinos2014high] to satisfy the I2M requirements for high resolution numerical simulations involving aneurysms. Despite the drawbacks of its Delaunay-based meshing techniques, PODM has been shown to generate good quality meshes for a segmented aneurysm image with reasonably good fidelity [EwCCBC3D]. The method also utilizes a parallel speculative execution model, exhibiting excellent performance when generating large meshes (e.g., 1 billion elements) on a large shared memory architecture (achieving a speedup of approximately 123 when utilizing 144 cores [foteinos2014high]). Consequently, we extend the method’s function approximation capability (introduced in [CNFPaper] to specifically address nuclear femtography problems) to now generate an adaptive isotropic mesh for an aneurysm case. To guarantee quality and fidelity, PODM adheres to specific criteria when determining if elements should be adapted (i.e., split). The user can specify their own criteria through a sizing function that can be loaded as a shared library, which PODM will utilize while generating a mesh in parallel. For our first test case (the segmented image), we construct a sizing function in a similar manner as described in section 4.1, based on [ComparisonPhysicsBased2023]. Algorithm 3 shows the sizing function that is utilized to determine if an element should be adapted. 3x3 positive definite metric tensor matrices are calculated for each of the element’s points based on their distance to the closest landmark [ComparisonPhysicsBased2023] (i.e., the closest point along the medial axis). Next, the sizing function checks if any of the element’s edges have a ”metric length” that is larger than a specified threshold. The metric length is a measure of how long an edge is in the space defined by the metric tensor [PLGeorgeMeshGeneration2008]. The average of the tensors at the two endpoints of an edge is calculated to obtain an interpolated tensor for that edge. If the calculated metric length exceeds the threshold, the element is marked for refinement. PODM’s output mesh quality and performance when it uses this sizing function for volume adaptivity (and when it does not, i.e., only boundary adaptivity based on the curvature) is evaluated in section 5.
Algorithm 3 PODM’s Function Approximation for an Aneurysm
SHOULD_BE_ADAPTED($P_{1}...P_{N}$, $L_{1}...L_{K}$, adapt_const)
Input: $P_{1}...P_{N}$ is the list of an element’s $N$ points
$L_{1}...L_{K}$ is the list of landmark points that affect adaptivity of elements throughout the mesh
adapt_const is a heuristic value that controls how aggressive adaptation should be
Output: Returns either true or false, designating if the element should be adapted
[…]
2:foreach p $\in$ $P_{1}...P_{N}$ do
3:  dist = distance between p and the closest landmark point among $L_{1}...L_{K}$
4:  diag = 1 / (dist * dist)
5:  $MT_{p},_{1}$ = diag
6:  $MT_{p},_{5}$ = diag
7:  $MT_{p},_{9}$ = diag
8:end for
9:edges = list of edge combinations among the points in $P_{1}...P_{N}$
10:foreach edge $\in$ edges do
11:  $edge\_point_{1}$ = first point in edge
12:  $edge\_point_{2}$ = second point in edge
13:  edge_vector.x = $edge\_point_{2}$.x - $edge\_point_{1}$.x
14:  edge_vector.y = $edge\_point_{2}$.y - $edge\_point_{1}$.y
15:  edge_vector.z = $edge\_point_{2}$.z - $edge\_point_{1}$.z
16:  foreach i $\in$ 1…9 do
17:   $avgTensor_{i}$ = ($MT_{edge\_point_{1},_{i}}$ + $MT_{edge\_point_{2},_{i}}$) / 2
18:  end for
19:  metric_length = $\sqrt{GET\_METRIC\_LENGTH\_SQUARED(edge\_vector,avgTensor)}$
20:  if metric_length ¿ adapt_const then
21:   return true
22:  end if
23:end for
24:return false
GET_METRIC_LENGTH_SQUARED(edge_vector,metric_tensor)GET\_METRIC\_LENGTH\_SQUARED(edge\_vector,metric\_tensor)
1:return edge_vector.xedge\_vector.x * ($metric\_tensor_{1}$ * edge_vector.xedge\_vector.x + $metric\_tensor_{2}$ * edge_vector.yedge\_vector.y + $metric\_tensor_{3}$ * edge_vector.zedge\_vector.z) + edge_vector.yedge\_vector.y * ($metric\_tensor_{4}$ * edge_vector.xedge\_vector.x + $metric\_tensor_{5}$ * edge_vector.yedge\_vector.y + $metric\_tensor_{6}$ * edge_vector.zedge\_vector.z) + edge_vector.zedge\_vector.z * ($metric\_tensor_{7}$ * edge_vector.xedge\_vector.x + $metric\_tensor_{8}$ * edge_vector.yedge\_vector.y + $metric\_tensor_{9}$ * edge_vector.zedge\_vector.z)
5 Results
We first describe how our experiment is set up in terms of the test cases and the supercomputers utilized to test the proposed methods in section 5.1. We then evaluate the adaptive anisotropic I2M pipeline and adaptive isotropic I2M method in section 5.2, followed by a study of the hierarchical load balancing model in section 5.3 and the optimized local reconnection algorithm when generating large meshes in section 5.4.
5.1 Experimental Setup
The first aneurysm case is based on a rotational angiography scan of a carotid cavernous aneurysm with image spacing of 1.00 x 1.00 x 1.00 mm3 and an image size of 512 x 512 x 508 voxels3. The CBC3D mesh generated from it can be seen in Figure LABEL:fig:artery1, in addition to its approximate medial axis. The second case is a surface mesh of a middle cerebral artery bifurcation aneurysm given as a Piecewise Linear Complex (PLC) in a .stl format (seen in Figure LABEL:fig:artery2_plc_2). Figure 5(a) shows the vascular structure imaged by rotational angiography of the right carotid artery for the second aneurysm case, in addition to the given surface mesh obtained from a segmentation of the region of interest. Figure 5(b) shows streamline data from a CFD simulation for this case (executed outside of this study [CFDDrivenTopology2024]). Recall that we evaluate the adaptive anisotropic mesh generation method with metric tensor fields that are constructed using different techniques for each case. The full pipeline of the proposed approach is tested with the first case while some of the methods attempt to perform mesh generation/adaptation using only the surface mesh for the second case. All codes were compiled using GNU GCC 11.4.1. Data were collected on three supercomputers - Purdue University’s Anvil [AnvilCommunity2014, Anvil2022] and Old Dominion University’s Wahab and Turing supercomputers. On Anvil, we utilized a dual socket node that features two AMD EPYC 7763 CPUs @ 2.45 GHz (64 slots each) and 256 GB of memory. On Wahab, we utilized a dual socket node that features two Intel Xeon Gold 6148 CPUs @ 2.40 GHz (20 slots each) and 384GB of memory. Finally, a dual socket node in the Turing supercomputer features two Intel Xeon E5-2698 v3 CPUs @ 2.30 GHz (16 slots each) and 128 GB of memory. For a performance evaluation, each run in the following experiments was executed three times and the results were averaged.
(a) The original image of the second (middle cerebral artery bifurcation) aneurysm case is shown from an anteroposterior view. (a) shows the vascular structure imaged by rotational angiography of the right carotid artery. (b) zooms in to the region of interest while (c) highlights the PLC obtained from the segmentation of the aneurysm. (d) shows only the resulting PLC.
(b) Shown is streamline data of the velocity field from a CFD simulation with the second (middle cerebral artery bifurcation) aneurysm case. (a) and (b) show different viewpoints. The velocity color legend is in m/s [CFDDrivenTopology2024].
For the generated isotropic meshes, qualitative results are examined with respect to element dihedral angles while quality is examined with respect to metric conformity for the adapted anisotropic meshes. The goal of metric conformity is to create a unit grid, where edges are unit-length and elements are unit-volume with respect to the target metric. For calculating edge length and element mean ratio, we adopted the same definitions that appear in [ALAUZETSizeGradation]. Given that optimal edges should be unit-length, edges with length above or below one are considered to be sub-optimal. The measure for mean ratio is bounded between zero and one since it is normalized by the volume of an equilateral element. One is the optimal quality for an element’s mean ratio shape.
We also evaluate the performance of our parallel adaptive anisotropic mesh generation method by scaling the complexity of the metric tensor fields to generate much larger meshes (up to 100 million elements). The complexity C of a continuous metric field $\mathcal{M}$ is defined as:
$C(\mathcal{M})=\int_{\Omega}\sqrt{det(\mathcal{M}(x))}dx.$ (1)
Complexity on the discrete grid is computed by sampling $\mathcal{M}$ at each vertex i as the discrete metric field M,
$C(M)=\sum_{i=1}^{N}\sqrt{det(M_{i})}V_{i},$ (2)
where $V_{i}$ is the volume of the Voronoi dual surrounding each node. The complexity of a grid is known to have a linear dependency with respect to the number of points and tetrahedra, shown theoretically in [LoseilleAlauzetCMFI] and experimentally verified in [LoseilleAlauzetCMFII, Park2015ComparingAO]. The number of vertices are approximately 2C while the number of tetrahedra are approximately 12C. As shown in [LoseilleAlauzetCMFI, TsolakisEvaluation], scaling the complexity of a metric can generate the same relative distribution of element density and shape over a uniformly refined grid compared to the original complexity. The metric tensor $M_{C_{r}}$ that corresponds to the target complexity $C_{r}$ is evaluated by [LoseilleAlauzetCMFI]:
$M_{C_{r}}=\left(\frac{C_{r}}{C(M)}\right)^{\frac{2}{3}}M,$ (3)
where M is the metric tensor before scaling and C(M) is the complexity of the discrete metric before scaling.
5.2 Quantitative and Qualitative Analysis of I2M Pipeline
@@
The results obtained from CDT3D are also compared to PODM [foteinos2014high] and TetGen [TetGen2015] (replacing the adaptive anisotropic method with these adaptive isotropic methods). TetGen is a sequential, Delaunay-based method that generates isotropic tetrahedral meshes. In addition to an anisotropic metric field, an isotropic sizing function is also constructed (based on the aforementioned approach in [ComparisonPhysicsBased2023]) specifically for PODM (also shown in algorithm 3) and TetGen. We compare results between meshes generated by PODM and TetGen when they utilize this sizing function and when they do not (PODM’s default image-to-mesh conversion for the first case and TetGen’s tetrahedralization of the CBC3D surface). Additionally, we compare the PODM and TetGen isotropic meshes to isotropic meshes generated by CDT3D (again simply processing the same surface as a PLC - .stl).
@@
Table 1 shows the parameters used for each mesh generation method in the pipeline (except CBC3D). The first aneurysm case was also used to evaluate CBC3D in [EwCCBC3D] (as it was compared to other isotropic I2M conversion methods from industry and academia). The same parameters utilized for CBC3D in that study are utilized in this evaluation. A wide range of values were tested for the value of k used in the k-nn search when building the sizing function and metric tensor fields for the first case. The optimal parameters were determined based on the best quality mesh that was generated. If certain parameters are not listed in Table 1, then their default value was utilized. The optimal value for k with the isotropic runs of TetGen and PODM is 30. The optimal value for k in the metric-adapted anisotropic runs of CDT3D is 40. The CDT3D parameters nqual and nsmth were both increased to 10 while cdfn was changed to 0.8 to improve the final anisotropic mesh quality for the second aneurysm case (with no surface adaptation by CDT3D, explained later).
@@
TetGen (w/ adaptivity) r reconstruct mesh
@@
PODM $\delta$/1.2 element size
nthreads/* parallel threads
adapt_const/2 heuristic for level of adaptation
@@
CBC3D takes approximately 61.59 seconds (1 min) to convert the segmented image of the first aneurysm into a mesh of approximately 272K tetrahedra and 63K points. Figures 9(a) and 9(b) show cross sections of the volume meshes generated by each method, merged with the conforming AFLR boundary layer mesh. Figure 9(c) shows cross sections of the volume meshes generated for the second aneurysm case. Table 2 shows the number of elements generated by each method when processing the CBC3D mesh of aneurysm 1 and the surface mesh of aneurysm 2. The TetGen meshes contain fewer elements compared to the CDT3D meshes in both cases. Tables 3 and 4 show the runtime of each method for aneurysm cases 1 and 2, respectively. They also show the mesh generation/adaptation rates of each method (measured in elements generated per second). TetGen exhibits good performance when generating its small-size meshes. Although it is not listed in Table 3, it should be noted that CDT3D exhibits good scalability, as its runtime is reduced to approximately 1 minute and its anisotropic adaptation rate becomes 27K when utilizing 40 cores. When utilizing 10 cores, CDT3D’s mesh generation/adaptation rate exceeds that of TetGen’s. For the second case, CDT3D completes anisotropic adaptation in about 49 seconds when utilizing 40 cores. Using the same number of cores, PODM takes about 2 seconds (about 260K elements/sec) to generate an isotropic mesh for the first case without utilizing the sizing function and 4 seconds (about 135K elements/sec) when it does.
(a) Cross sections are shown of the volume meshes generated by CBC3D and TetGen for the first aneurysm case, merged with the conforming boundary layer mesh generated by AFLR.
(b) Cross sections are shown of the volume meshes generated by PODM and CDT3D for the first aneurysm case, where the CDT3D meshes are merged with the conforming boundary layer mesh generated by AFLR.
(c) Cross sections are shown of the volume meshes generated by each method for the second aneurysm case, merged with the conforming boundary layer mesh generated by AFLR.
Table 2: Number of tetrahedra and points generated by each method when processing either the segmented image (PODM) or the CBC3D mesh converted from the segmented image of the first (carotid cavernous) aneurysm and the surface mesh of the second (middle cerebral artery bifurcation) aneurysm. ‘Tets’ is short for tetrahedra. K means thousand and M means million.
@@
TetGen (w/ adaptivity) Tets 120K -
@@
PODM (no vol. adaptivity) Tets 599K -
Points 131K -
PODM (w/ vol. adaptivity) Tets 644K -
Points 141K -
AFLR (bl) Tets 912K 82K
Points 183K 16K
@@
Table 3: The time spent (approximately in minutes) and the number of elements generated per second by each method (in parentheses) for small-size meshes on the Wahab supercomputer is shown when processing either the segmented image (PODM) or the CBC3D mesh converted from the segmented image of the first (carotid cavernous) aneurysm. Instead of elements per second, the values in parentheses for Build_Sizing/Metric represent the CBC3D volume mesh points processed per second. ‘K’ means thousand.
@@
TetGen (w/ adaptivity) 0.1 (20K) - -
TetGen (no adaptivity) 0.06 (32K) - -
PODM (no vol. adaptivity) 0.35 (28K) 0.05 (167K) 0.04 (217K)
PODM (w/ vol. adaptivity) 1.1 (9K) 0.15 (67K) 0.1 (101K)
AFLR (bl) 0.43 (35K) - -
@@
Table 4: The time spent (approximately in seconds) and the number of elements generated per second by each method (in parentheses) for small-size meshes on the Wahab supercomputer is shown when processing the surface mesh of the second (middle cerebral artery bifurcation) aneurysm. ‘K’ means thousand.
@@
Mesh fidelity is evaluated using a two-sided Hausdorff Distance (HD) metric (based on an open source implementation [Commandeur]): $\text{HD}=\max\{\text{HD}_{\text{I}\rightarrow\text{M}},\text{HD}_{\text{M}\rightarrow\text{I}}\}$, where $\text{HD}_{\text{I}\rightarrow\text{M}}$ is the value of the metric from the image to the mesh, and $\text{HD}_{\text{M}\rightarrow\text{I}}$ is the value of the metric from the mesh to the image. A low HD error indicates a high fidelity. Hausdorff Distance is computed between two point sets. The first point set contains the vertices located on the mesh surface and the second point set contains the voxels located on the boundaries of the segmented material. The HD for the CBC3D-generated mesh of the first aneurysm case (in Figure LABEL:fig:artery1) is 4.51. The meshes generated by PODM (Figures LABEL:fig:podm_no_adapt_artery1_cut and LABEL:fig:podm_adapt_artery1_cut) also exhibit good fidelity, as their HD are both 2.42. The fidelity of the meshes generated by CBC3D and PODM were also evaluated for the first aneurysm case in [EwCCBC3D]. They were both compared to several highly used I2M methods, where CBC3D exhibited good fidelity while maintaining better smoothness compared to the other methods’ meshes (including PODM). Irrespective of PODM, surface refinement/adaptation is turned off for all mesh generation methods tested in the pipeline (except CBC3D). This ensures that the high fidelity and smooth surface generated by CBC3D remains intact.